AB INITIO STUDY OF THE AMERICYL AND CURYL IONS

\underline{SCOTT R. BROZELL} and RUSSELL M. PITZER, Department of Chemistry, The Ohio State University, Columbus, OH 43210.

Restricted Hartree-Fock (RHF) and spin-orbit configuration-interaction (SOCI) calculations were performed on the ground and low-lying excited states of the americyl ion, AmO_{2}^{2+}. The RHF δ_{u}^{2} ϕ_{u}^{1} ,^{4}Φ_{u} ground state symmetric stretch potential energy curve yields a frequency of 900 cm^{−1} compared to an experimental value of 730 cm^{−1} from Raman spectroscopy on aqueous solutions^{a} of AmO_{2}^{1+}. The SOCI ground state is δ_{u}^{2} ϕ_{u}^{1} ,^{4}Φ_{3/2u}; the excited states are regular components from the above term followed by irregular components from the δ_{u}^{1} ϕ_{u}^{2} ,^{4}Δ_{u} term. The low energy transitions are f → f. Analogous calculations are planned for the curyl ion, CmO_{2}^{2+}. The actinides are modeled with relativistic effective core potentials and Gaussian correlation consistent double-zeta plus polarization (cc-pVDZ) basis sets.

aG. M. Begun et al., Inorg. Chem. 23, 1914 (1984).